A train 110m long passesa man, running at 6 kmph in the direction opposite to that ofthe train ,in 6 seconds.What is the speed of the train?

Grade:8

2 Answers

Arun
6 years ago
 DerD student x=60 km/hrx= 66-6therefore x+6 = 66let speed of train= x= 66 km/hrspeed of train relative to man: 110/6 * 18/5 km/hr  RegardsArun (askIITians forum expert)
Amit mondal
6 years ago
Let train speed be X km/hThen relative speed will be (X-6)convert relative speed into (X-6)×5/18 m/secrequired time = length of the train/ relative speed6= 110×18/(X-6)×5X= 72 m/sec